How much of a 22 percent alcohol plus a 9 percent alcohol do you need to get 17.5 percent alcohol content in 560 ml. Need formulas and answer. Thanks
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 22.00% ---------------- x ml
Alcohol II 9.00% ------ 560 - x ml
Mixture 17.50% ---------------- 560
Total 560 ml
22.00% x + 9.00% ( 560 - x ) = 17.50% * 560
22 x + 9 ( 560 - x ) = 9800
22 x + 5040 - 9 x = 9800
22 x - 9 x = 9800 - -5040
13 x = 4760
/ 13
x = 366.15 ml 22.00% Alcohol I
193.85 ml 9.00% Alcohol II

m.ananth@hotmail.ca